# Batayan: A Filipino NLP benchmark for evaluating Large Language Models

### v1

@@ Appendix A Inter-rater Agreement
Table 4: Joint agreement is calculated as the percentage of times all the raters uninamously agree that the samples fulfill the criteria. Rating is calculated as the average score from 0 to 3 assigned by the raters to the samples under the given criteria.
Appendix B Overview of Tasks and Datasets
Competency Task Dataset License
NLU PI PAWS Zhang et al. (2019) CC BY 4.0
QA Belebele Bandarkar et al. (2024) CC BY-NC 4.0
SA PH Elections Cabasag et al. (2019) Unknown
TD PH Elections Cabasag et al. (2019) Unknown
NLR CR Balanced COPA Kavumba et al. (2019) CC BY 4.0
NLI XNLI Conneau et al. (2018) CC BY-NC 4.0
NLG AS XL-Sum Hasan et al. (2021) CC BY-NC-SA 4.0
MT FLORES 200 NLLB Team et al. (2022) CC BY-SA 4.0
Table 5: License details for datasets used in Batayan.
Competency Task Label No. Samples
NLU PI True 200
False 200
SA Negative (Negatibo) 200
Neutral (Neutral) 200
Positive (Positibo) 200
TD Clean (Malinis) 200
Toxic (Mapoot) 200
NLR CR Cause (Sanhi) 200
Effect (Bunga) 200
NLI Contradiction 200
Entailment 200
Neutral 200
Table 6: Class distribution per task.
Competency Task Total No. Samples Avg. No. Words/Sample
NLU PI 400 22.50 (Sentence 1), 22.47 (Sentence 2)
QA 100 14.75 (Question), 4.26 (Choices)
SA 600 24.32
TD 400 20.73
NLR CR 400 7.24 (Premise), 6.05 (Choices)
NLI 600 23.30 (Sentence 1), 12.20 (Sentence 2)
NLG AS 100 120.8
MT (English Text) 600 21.42
MT (Tagalog Text) 600 25.05
Table 6 shows the distribution of class labels for tasks that involve classification. We On the other hand, Table 7 presents quantitative statistics on dataset size and average word counts. The variation in text length reflects task-specific requirements: AS involves longer passages, while classification tasks (SA, TD, and PI) typically contain more concise text samples. The QA dataset, in particular, has relatively short questions and answer choices, mirroring real-world multiple-choice assessments. Further granularity is provided in Table 8, which breaks down sentence, word, and character-level statistics. This statistics shows structural differences between tasks; for instance, MT has sentence-level parallelism, while CR and NLI involve distinct premise-hypothesis or question-choice relationships.
NLU PI Sentence 1 1.0 22.50 123.01
PI Sentence 2 1.0 22.47 122.99
QA Question 1.0 14.75 80.62
QA Choice 1 1.0 4.66 29.10
QA Choice 2 1.0 4.61 28.60
QA Choice 3 1.0 4.76 29.74
QA Choice 4 1.0 3.0 28.58
SA Text 2.21 24.32 120.44
TD Text 2.45 20.73 99.37
NLR CR Premise 1.0 7.24 37.55
CR Choice 1 1.0 6.05 30.91
CR Choice 2 1.0 6.05 30.62
NLI Sentence 1 1.0 23.30 127.88
NLI Sentence 2 1.0 12.20 67.04
NLG AS Text 5.43 120.8 666.18
MT English Text 1.12 21.42 113.07
MT Filipino Text 1.12 25.05 140.58
Appendix C Prompt Templates
Table 9: Example instances for each task in Batayan.
@@ Appendix C Prompt Templates
Appendix D Experimental Setup
The default evaluation setting for Batayan is zero-shot prompting, and performance scores are calculated based on a single run. Despite being a comprehensive benchmark, Batayan requires very few resources to execute. For our experiments, we allocated a computational budget of 0.25 GPU hours on 1 H100 GPU per model, totaling to 2 H100 GPU hours per run of Batayan.
Table LABEL:tab:model_sizes reports the number of parameters in the models evaluated, while Table LABEL:tab:inference_parameters details the decoding parameters used in the experiments.
Models No. parameters
aisingapore/gemma2-9b-cpt-sea-lionv3-instruct 9B
aisingapore/llama3.1-8b-cpt-sea-lionv3-instruct 8B
CohereForAI/aya-expanse-8b 8B
google/gemma-2-9b-it 9B
meta-llama/Llama-3.1-8B-Instruct 8B
Qwen/Qwen2.5-7B-Instruct 7B
sail/Sailor2-8B-Chat 8B
SeaLLMs/SeaLLMs-v3-7B-Chat 7B
@@ Appendix D Experimental Setup
Appendix E Dataset Examples
Competency Task Example
Label: Paraprase (Paraphrase)
Question: Ano ang pinaniniwalaang dahilan ng pagbagsak?
Label: 1
Label: Negatibo (Negative)
Label: Malinis (Clean)
NLR CR Text: Hindi tinanggap ang tsekeng ginawa ko.
Question: Sanhi (Cause)
Label: 0
Sentence 2: Hindi kami pupunta sa birthday party ni Tita Basia ngayon.
Label: Contradiction

### v2

@@ Appendix C Inter-rater Agreement
Table 7: Joint agreement is calculated as the percentage of times all the raters unanimously agree that the samples fulfill the criteria. Rating is calculated as the average score assigned by the raters to the samples under the given Likert-scaled (0-3) criteria.
Appendix D Prompt Templates
Table 8: Performance of commercial LLMs on Batayan NLG tasks.
@@ Appendix D Prompt Templates
Appendix E Details of Models Evaluated
Models Source Link No. params
Base pre-trained models
aisingapore/Gemma-SEA-LION-v3-9B https://huggingface.co/aisingapore/Gemma-SEA-LION-v3-9B 9B
aisingapore/Llama-SEA-LION-v2-8B https://huggingface.co/aisingapore/Llama-SEA-LION-v2-8B 8B
aisingapore/Llama-SEA-LION-v3-8B https://huggingface.co/aisingapore/Llama-SEA-LION-v3-8B 8B
aisingapore/Llama-SEA-LION-v3-70B https://huggingface.co/aisingapore/Llama-SEA-LION-v3-70B 70B
google/gemma-2-9b https://huggingface.co/google/gemma-2-9b 9B
google/gemma-2-27b https://huggingface.co/google/gemma-2-27b 27B
google/gemma-3-12b-pt https://huggingface.co/google/gemma-3-12b-pt 12B
google/gemma-3-27b-pt https://huggingface.co/google/gemma-3-27b-pt 27B
meta-llama/Llama-3.1-8B https://huggingface.co/meta-llama/Llama-3.1-8B 8B
meta-llama/Llama-3.1-70B https://huggingface.co/meta-llama/Llama-3.1-70B 70B
Qwen/Qwen2-7B https://huggingface.co/Qwen/Qwen2-7B 7B
Qwen/Qwen2-72B https://huggingface.co/Qwen/Qwen2-72B 72B
Qwen/Qwen2.5-7B https://huggingface.co/Qwen/Qwen2.5-7B 7B
Qwen/Qwen2.5-14B https://huggingface.co/Qwen/Qwen2.5-14B 14B
Qwen/Qwen2.5-32B https://huggingface.co/Qwen/Qwen2.5-32B 32B
Qwen/Qwen2.5-72B https://huggingface.co/Qwen/Qwen2.5-72B 72B
sail/Sailor2-8B https://huggingface.co/sail/Sailor2-8B 8B
sail/Sailor2-20B https://huggingface.co/sail/Sailor2-20B 20B
SeaLLMs/SeaLLMs-v3-7B https://huggingface.co/SeaLLMs/SeaLLMs-v3-7B 7B
Instruction-tuned models
aisingapore/Gemma-SEA-LION-v3-9B-IT https://huggingface.co/aisingapore/Gemma-SEA-LION-v3-9B-IT 9B
aisingapore/Llama-SEA-LION-v2-8B-IT https://huggingface.co/aisingapore/Llama-SEA-LION-v2-8B-IT 8B
aisingapore/Llama-SEA-LION-v3-8B-IT https://huggingface.co/aisingapore/Llama-SEA-LION-v3-8B-IT 8B
aisingapore/Llama-SEA-LION-v3-70B-IT https://huggingface.co/aisingapore/Llama-SEA-LION-v3-70B-IT 70B
CohereForAI/aya-23-8B https://huggingface.co/CohereForAI/aya-23-8B 8B
CohereForAI/aya-23-35B https://huggingface.co/CohereForAI/aya-23-35B 35B
CohereForAI/aya-expanse-8b https://huggingface.co/CohereForAI/aya-expanse-8b 8B
CohereForAI/aya-expanse-32b https://huggingface.co/CohereForAI/aya-expanse-32b 32B
deepseek-ai/DeepSeek-V3 https://huggingface.co/deepseek-ai/DeepSeek-V3 671B
google/gemma-2-9b-it https://huggingface.co/google/gemma-2-9b-it 9B
google/gemma-2-27b-it https://huggingface.co/google/gemma-2-27b-it 27B
google/gemma-3-12b-it https://huggingface.co/google/gemma-3-12b-it 12B
google/gemma-3-27b-it https://huggingface.co/google/gemma-3-27b-it 27B
MERaLiON/Llama-3-MERaLiON-8B-Instruct https://huggingface.co/MERaLiON/Llama-3-MERaLiON-8B-Instruct 8B
meta-llama/Llama-3.1-8B-Instruct https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct 8B
meta-llama/Llama-3.1-70B-Instruct https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct 70B
meta-llama/Llama-3.3-70B-Instruct https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct 70B
Qwen/Qwen2-7B-Instruct https://huggingface.co/Qwen/Qwen2-7B-Instruct 7B
Qwen/Qwen2-72B-Instruct https://huggingface.co/Qwen/Qwen2-72B-Instruct 72B
Qwen/Qwen2.5-7B-Instruct https://huggingface.co/Qwen/Qwen2.5-7B-Instruct 7B
Qwen/Qwen2.5-14B-Instruct https://huggingface.co/Qwen/Qwen2.5-14B-Instruct 14B
Qwen/Qwen2.5-32B-Instruct https://huggingface.co/Qwen/Qwen2.5-32B-Instruct 32B
Qwen/Qwen2.5-72B-Instruct https://huggingface.co/Qwen/Qwen2.5-72B-Instruct 72B
sail/Sailor2-8B-Chat https://huggingface.co/sail/Sailor2-8B-Chat 8B
sail/Sailor2-20B-Chat https://huggingface.co/sail/Sailor2-20B-Chat 20B
SeaLLMs/SeaLLMs-v3-7B-Chat https://huggingface.co/SeaLLMs/SeaLLMs-v3-7B-Chat 7B
Tower-Babel/Babel-9B-Chat https://huggingface.co/Tower-Babel/Babel-9B-Chat 9B
Tower-Babel/Babel-83B-Chat https://huggingface.co/Tower-Babel/Babel-83B-Chat 83B
Reasoning models
deepseek-ai/DeepSeek-R1 https://huggingface.co/deepseek-ai/DeepSeek-R1 671B
Commercial models
google/gemini-1.5-flash-002 https://cloud.google.com/vertex-ai/generative-ai/docs/learn/model-versions -
google/gemini-1.5-pro-002 https://cloud.google.com/vertex-ai/generative-ai/docs/learn/model-versions -
google/gemini-2.0-flash-001 https://cloud.google.com/vertex-ai/generative-ai/docs/learn/model-versions -
google/gemini-2.0-flash-lite-001 https://cloud.google.com/vertex-ai/generative-ai/docs/learn/model-versions -
openai/gpt-4o-mini-2024-07-18 https://platform.openai.com/docs/models/gpt-4o-mini -
openai/gpt-4o-2024-08-06 https://platform.openai.com/docs/models/gpt-4o -
Model Reference Filipino IT Rationale for selection
Aya 23 and Expanse Aryabumi et al. (2024); Dang et al. (2024) LLMs with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
Babel Zhao et al. (2025) LLMs with multilingual support including Filipino. This model was pre-trained and fine-tuned on Filipino tokens.
DeepSeek V3 and R1 DeepSeek-AI (2025) LLMs with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
Gemma 2 and 3 Team et al. (2024b, 2025) LLMs with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
Gemini 1.5 and 2 Team et al. (2024a) Commercial AI systems with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
GPT-4o and 4o-mini OpenAI et al. (2024) Commercial AI systems with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
Llama 3 Grattafiori et al. (2024) LLMs with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
MERaLiON Huang et al. (2025) LLM with support for English, Chinese, and Indonesian. Llama 3 model was continue pre-trained on 120B English, Chinese, and Indonesian tokens. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
Qwen 2 and 2.5 Yang et al. (2024); Qwen et al. (2025) LLMs with multilingual support, including Tagalog.
Sailor 2 Dou et al. (2025) LLMs with support for Southeast Asian languages. Qwen 2.5 7B model was continue pre-trained on 400B SEA tokens, and was specifically fine-tuned on Filipino instructions.
SEA-LION v2 and v3 Ng et al. (2025) LLMs with support for Southeast Asian languages. Llama 3 and Gemma 2 series models were continue pre-trained on 200B tokens across 11 Southeast Asian languages, and was specifically fine-tuned on Filipino instructions.
SeaLLMs v3 Zhang et al. (2024) LLMs with support for Southeast Asian languages. These models were pre-trained and fine-tuned on Filipino tokens.
Appendix F Experimental Setup
The default evaluation setting for Batayan is zero-shot prompting, however base pre-trained models with no instruction-tuning are evaluated on a five-shot setting. Performance scores are calculated based on a single run. Table LABEL:tab:inference_parameters details the decoding parameters used in the experiments.
@@ Appendix F Experimental Setup
Appendix G Complete Experimental Results
Model PI QA SA TD NLU Avg. CR NLI NLR Avg.
Base pre-trained models (five-shot)
aisingapore/Gemma-SEA-LION-v3-9B 82.40 89.02 67.60 51.54 72.64 0.00a 63.93 31.97
aisingapore/Llama-SEA-LION-v2-8B 71.21 69.54 48.86 52.89 60.62 0.00a 28.99 14.49
aisingapore/Llama-SEA-LION-v3-8B 71.96 71.07 38.58 66.44 62.01 0.00a 48.59 24.30
google/gemma-2-9b 70.64 85.79 51.09 74.26 70.44 0.00a 46.04 23.02
meta-llama/Llama-3.1-8B 60.31 71.91 50.16 58.51 60.22 0.00a 26.82 13.41
Qwen/Qwen2-7B 84.69 74.12 72.11 65.48 74.10 0.00a 45.85 22.92
Qwen/Qwen2.5-7B 84.48 69.89 72.22 64.71 72.82 0.00a 41.11 20.55
sail/Sailor2-8B 79.44 75.93 70.25 67.64 73.32 0.00a 43.84 21.92
SeaLLMs/SeaLLMs-v3-7B 84.95 74.05 71.12 63.54 73.41 0.00a 41.46 20.73
Instruction-tuned models (zero-shot)
aisingapore/Gemma-SEA-LION-v3-9B-IT 82.00 82.80 75.99 72.85 78.41 92.75 68.64 80.69
aisingapore/Llama-SEA-LION-v2-8B-IT 68.50 45.41 49.81 58.04 55.44 50.23 37.63 43.93
aisingapore/Llama-SEA-LION-v3-8B-IT 74.97 80.69 51.24 60.57 66.87 79.44 65.87 72.65
CohereForAI/aya-23-8B 24.68 33.50 33.49 32.70 31.09 28.12 16.67 22.39
CohereForAI/aya-expanse-8b 32.65 2.97a 57.44 34.39 31.86 1.96a 25.27 13.61
google/gemma-2-9b-it 82.74 83.96 70.80 64.48 75.50 91.25 60.24 75.74
MERaLiON/Llama-3-MERaLiON-8B-Instruct 23.47 64.90 35.16 27.64 37.79 28.97 23.35 26.16
meta-llama/Llama-3.1-8B-Instruct 61.38 74.90 52.55 50.00 59.71 72.66 56.28 64.47
Qwen/Qwen2-7B-Instruct 73.08 56.01 52.63 26.10 51.96 43.99 32.35 38.17
Qwen/Qwen2.5-7B-Instruct 85.22 69.07 67.82 60.84 70.74 46.15 57.80 51.98
sail/Sailor2-8B-Chat 73.60 59.48 48.67 47.24 57.25 57.52 61.21 59.36
SeaLLMs/SeaLLMs-v3-7B-Chat 37.85 49.72 70.63 61.48 54.92 71.82 20.72 46.27
Tower-Babel/Babel-9B-Chat 78.76 61.68 72.52 56.74 67.42 83.00 26.89 54.95
  • a
    Could not follow answer instructions.
Model PI QA SA TD NLU Avg. CR NLI NLR Avg.
Base pre-trained models (five-shot)
google/gemma-2-27b 72.63 86.04 69.60 75.95 76.05 0.00a 56.21 28.11
google/gemma-3-12b-pt 71.42 86.02 79.30 51.60 72.08 0.00a 59.47 29.74
google/gemma-3-27b-pt 59.29 84.83 69.84 52.34 66.58 0.00a 55.13 27.56
Qwen/Qwen2.5-14B 89.25 82.03 68.36 62.89 75.63 0.00a 53.10 26.55
Qwen/Qwen2.5-32B 87.69 84.98 67.71 66.62 76.75 0.00a 55.09 27.54
sail/Sailor2-20B 85.46 84.94 68.47 48.83 71.92 0.00a 63.09 31.55
Instruction-tuned models (zero-shot)
CohereForAI/aya-expanse-32b 76.39 54.82 63.25 74.70 67.29 55.46 66.57 61.01
google/gemma-3-12b-it 82.27 83.94 62.10 80.50 77.20 93.75 70.64 82.19
google/gemma-2-27b-it 70.68 69.18 70.39 80.36 72.65 67.76 80.25 73.04
google/gemma-3-27b-it 70.20 87.91 67.77 81.74 76.91 94.75 67.73 81.24
Qwen/Qwen2.5-14B-Instruct 63.75 64.64 72.34 72.73 68.37 53.47 43.83 48.65
Qwen/Qwen2.5-32B-Instruct 83.70 84.97 74.60 75.50 79.69 85.98 60.58 73.28
sail/Sailor2-20B-Chat 39.17 13.81 54.41 74.03 45.36 22.98 47.88 35.43
  • a
    Could not follow answer instructions.
Model PI QA SA TD NLU Avg. CR NLI NLR Avg.
Base pre-trained models (five-shot)
aisingapore/Llama-SEA-LION-v3-70B 83.48 87.02 45.89 77.49 73.47 0.00a 79.62 39.81
meta-llama/Llama-3.1-70B 83.49 86.05 59.72 51.70 70.24 0.00a 55.78 27.89
Qwen/Qwen2-72B 90.25 88.86 67.88 77.51 81.13 0.00a 69.63 34.81
Qwen/Qwen2.5-72B 86.89 86.79 74.73 79.61 82.01 0.00a 67.58 33.79
Instruction-tuned models (zero-shot)
aisingapore/Llama-SEA-LION-v3-70B-IT 84.07 87.01 68.79 76.99 79.21 94.75 75.34 85.04
CohereForAI/aya-23-35B 81.83 61.46 61.68 53.06 64.51 74.16 40.54 57.35
deepseek-ai/DeepSeek-V3 83.07 90.00 81.78 76.21 82.77 94.25 63.25 78.75
meta-llama/Llama-3.1-70B-Instruct 83.57 88.13 67.29 63.80 75.70 93.49 72.62 83.06
meta-llama/Llama-3.3-70B-Instruct 82.62 86.96 66.12 77.40 78.28 94.25 72.97 83.61
Qwen/Qwen2-72B-Instruct 85.37 86.10 76.16 76.38 81.00 58.54 68.39 63.47
Qwen/Qwen2.5-72B-Instruct 80.40 84.00 75.32 73.44 78.29 91.75 63.95 77.85
Tower-Babel/Babel-83B-Chat 85.06 33.12 73.20 81.24 68.15 49.62 68.14 58.88
Reasoning models (zero-shot)
deepseek-ai/DeepSeek-R1 79.84 87.93 71.47 71.47 77.68 97.00 71.06 84.03
  • a
    Could not follow answer instructions.
Model PI QA SA TD NLU Avg. CR NLI NLR Avg.
google/gemini-1.5-flash-002 85.71 84.98 63.81 66.04 75.14 94.24 69.19 81.72
google/gemini-1.5-pro-002 85.62 89.95 69.48 79.27 81.08 95.74 76.71 86.23
google/gemini-2.0-flash-001 83.98 88.05 73.93 80.49 81.61 97.74 70.47 84.11
google/gemini-2.0-flash-lite-001 85.02 85.97 64.33 78.94 78.57 96.49 70.20 83.35
openai/gpt-4o-mini-2024-07-18 86.11 85.04 70.77 70.77 78.17 92.99 68.18 80.59
openai/gpt-4o-2024-08-06 88.21 71.88 74.78 74.78 77.41 97.25 73.33 85.29
Models AS MT (engtgl) MT (tgleng) NLG
BERTScore ChrF++ ROUGE-L ChrF++ MetricX-24 ChrF++ MetricX-24 Avg.
Base pre-trained models (five-shot)
aisingapore/Gemma-SEA-LION-v3-9B 77.01 34.19 33.71 58.90 82.75 56.63 91.15 62.05
aisingapore/Llama-SEA-LION-v2-8B 75.86 27.17 30.88 48.45 62.74 58.17 87.10 55.77
aisingapore/Llama-SEA-LION-v3-8B 75.73 32.41 30.56 53.10 72.41 57.70 88.91 58.69
google/gemma-2-9b 76.60 31.72 31.69 59.06 82.01 63.27 91.16 62.22
meta-llama/Llama-3.1-8B 75.80 29.93 30.13 51.94 71.10 58.44 88.77 58.01
Qwen/Qwen2-7B 73.84 27.64 25.71 42.61 47.37 53.60 82.13 50.41
Qwen/Qwen2.5-7B 73.26 26.71 25.04 38.89 40.28 42.59 81.58 46.91
sail/Sailor2-8B 77.05 30.69 34.50 62.01 86.30 63.41 91.05 63.57
SeaLLMs/SeaLLMs-v3-7B 73.78 28.37 26.57 43.45 47.79 44.54 83.74 49.75
Instruction-tuned models (zero-shot)
aisingapore/Gemma-SEA-LION-v3-9B-IT 73.00 34.16 20.99 57.79 87.24 58.63 90.56 60.35
aisingapore/Llama-SEA-LION-v2-8B-IT 73.39 32.91 24.14 50.90 64.62 48.49 55.04 49.93
aisingapore/Llama-SEA-LION-v3-8B-IT 73.34 33.80 23.43 57.14 82.67 58.33 87.69 59.49
CohereForAI/aya-23-8B 67.37 24.01 11.76 15.61 9.68 46.20 58.59 33.32
CohereForAI/aya-expanse-8b 59.48 6.69 1.64 39.30 37.47 49.81 78.88 39.04
google/gemma-2-9b-it 72.73 34.19 21.81 56.80 84.23 50.97 76.62 56.76
MERaLiON/Llama-3-MERaLiON-8B-Instruct 74.08 31.79 25.68 49.77 64.85 41.66 60.14 49.71
meta-llama/Llama-3.1-8B-Instruct 74.24 34.27 26.94 44.79 51.76 46.33 70.65 49.85
Qwen/Qwen2-7B-Instruct 69.51 26.09 13.58 42.22 42.86 45.13 50.01 41.34
Qwen/Qwen2.5-7B-Instruct 70.61 28.89 15.86 38.99 34.57 49.02 76.36 44.91
sail/Sailor2-8B-Chat 71.82 31.83 19.75 54.04 80.93 35.41 19.66 44.78
SeaLLMs/SeaLLMs-v3-7B-Chat 70.34 27.84 16.49 46.36 61.07 43.27 60.37 46.53
Tower-Babel/Babel-9B-Chat 71.61 31.55 18.93 49.97 72.50 48.69 75.69 52.70
Models AS MT (engtgl) MT (tgleng) NLG
BERTScore ChrF++ ROUGE-L ChrF++ MetricX-24 ChrF++ MetricX-24 Avg.
Base pre-trained models (five-shot)
google/gemma-2-27b 77.27 33.71 33.50 61.16 85.56 64.77 91.82 63.97
google/gemma-3-12b-pt 77.42 32.12 34.32 61.66 85.52 64.69 92.08 63.97
google/gemma-3-27b-pt 77.70 35.53 35.36 62.89 87.06 64.98 92.27 65.11
Qwen/Qwen2.5-14B 74.39 30.34 27.87 49.01 62.69 59.41 88.63 56.05
Qwen/Qwen2.5-32B 75.13 29.73 29.31 50.94 65.06 60.18 89.34 57.10
sail/Sailor2-20B 76.73 32.41 33.47 61.75 86.30 63.88 91.73 63.75
Instruction-tuned models (zero-shot)
CohereForAI/aya-expanse-32b 71.18 30.12 16.85 49.14 62.76 57.35 85.05 53.21
google/gemma-2-27b-it 73.41 35.11 23.21 59.51 88.19 61.87 91.68 61.86
google/gemma-3-12b-it 72.10 32.80 19.43 60.09 89.39 61.85 91.61 61.04
google/gemma-3-27b-it 72.25 33.26 19.55 61.02 90.71 62.09 92.18 61.58
Qwen/Qwen2.5-14B-Instruct 71.92 31.85 19.24 44.78 54.51 58.01 86.08 52.34
Qwen/Qwen2.5-32B-Instruct 72.47 33.43 21.22 47.47 59.47 58.77 86.90 54.25
sail/Sailor2-20B-Chat 71.47 31.20 18.41 49.45 72.51 49.52 72.79 52.19
Models AS MT (engtgl) MT (tgleng) NLG
BERTScore ChrF++ ROUGE-L ChrF++ MetricX-24 ChrF++ MetricX-24 Avg.
Base pre-trained models (five-shot)
aisingapore/Llama-SEA-LION-v3-70B 76.87 35.49 33.66 60.70 85.28 63.39 91.52 63.85
meta-llama/Llama-3.1-70B 76.80 33.98 33.54 60.20 84.32 65.80 91.85 63.79
Qwen/Qwen2-72B 76.81 34.15 33.24 54.75 74.84 65.13 90.87 61.40
Qwen/Qwen2.5-72B 76.41 34.12 31.96 54.69 74.92 64.50 90.85 61.06
Instruction-tuned models (zero-shot)
aisingapore/Llama-SEA-LION-v3-70B-IT 73.25 34.82 23.81 61.02 86.24 63.11 91.28 61.93
CohereForAI/aya-23-35B 72.81 29.94 22.08 46.00 56.79 49.54 72.45 49.94
deepseek-ai/DeepSeek-V3 73.19 35.16 22.99 61.61 89.21 63.06 92.28 62.50
meta-llama/Llama-3.1-70B-Instruct 74.76 35.56 27.94 60.32 84.73 62.78 91.20 62.47
meta-llama/Llama-3.3-70B-Instruct 72.90 34.03 22.63 58.98 83.05 61.24 90.72 60.51
Qwen/Qwen2-72B-Instruct 67.05 25.08 11.38 52.84 70.63 60.05 88.73 53.68
Qwen/Qwen2.5-72B-Instruct 71.62 31.52 18.02 52.01 72.30 64.33 90.02 57.12
Tower-Babel/Babel-83B-Chat 67.72 20.58 13.35 53.29 76.61 51.56 85.42 52.65
Reasoning models (zero-shot)
deepseek-ai/DeepSeek-R1 71.97 32.70 19.59 59.15 90.91 61.01 92.30 61.09
Models AS MT (engtgl) MT (tgleng) NLG
BERTScore ChrF++ ROUGE-L ChrF++ MetricX-24 ChrF++ MetricX-24 Avg.
google/gemini-1.5-flash-002 72.44 34.00 21.20 61.01 90.11 65.23 92.04 62.29
google/gemini-1.5-pro-002 73.62 36.16 24.00 61.07 90.90 64.14 91.30 63.03
google/gemini-2.0-flash-001 72.99 34.85 21.96 63.88 90.87 65.64 92.53 63.25
google/gemini-2.0-flash-lite-001 72.83 34.80 21.87 63.45 90.06 65.32 91.51 62.83
openai/gpt-4o-mini-2024-07-18 72.70 33.97 21.67 62.46 88.59 63.12 91.68 62.03
openai/gpt-4o-2024-08-06 72.83 34.36 21.55 63.86 90.19 64.89 92.51 62.88